If albert answered 26 questions correctly and earned a score of 65% what would the percentage be if he answered 30 questions correctly?
Answer by rfer(16322)

26/0.65=40
30/40=0.75=75%
-----------------
0.65/26=x/30
26x=19.5
x=19.5/26
x=0.75=75%